A)Find the inverse of the function f(x)=3+{2x-1 B) SOLVE log(x+2)+log 3 (x-2)=2 5. 6. vords e here to search ORI
Answer #1

soln (A f(x)= 3 +522-T We have to find inverse. let y = f(x) y = 3+ 3/2x-1 and y by ni Now replace a by y a=3+$2y | x-3 = Ny-